# Possibilistic Instrumental Variable Regression

### v1

@@ 5 Conclusion
In this paper, we propose a method for instrumental variable regression based on possibility theory. This allows for valid inference under some user-specified violations of the instrument exogeneity assumption, providing a very natural way to perform sensitivity analyses. The resulting inference is based on first principles and directly reflects the epistemic uncertainty about the parameters. Our proposed approach performs well in simulation experiments and delivers credible results in a real-data application.
@@ A.1 The structural posterior possibility function
$\begin{aligned} \log f_{\mathrm{S}}(\alpha,\beta,\Sigma\mid W)&=-\frac{n}{2}\log\lvert R(\beta)\Sigma R(\beta)^{\intercal}\rvert\\ &-\frac{1}{2}\text{tr}\left((R(\beta)\Sigma R(\beta)^{\intercal})^{-1}W^{\intercal}M_{Z}W\right)\\ &-\frac{1}{2}(\alpha-t(\beta))^{\intercal}\frac{Z^{\intercal}Z}{\sigma_{11}}(\alpha-t(\beta)).\end{aligned}$ (5)
@@ A.2 Proof of Proposition 1 and computational considerations
$f(\beta\mid\alpha\in A,W)=\sup_{\alpha\in A,\,\Sigma\in\mathbb{S}^{2}_{+}}\frac{f_{\mathrm{S}}(\alpha,\beta,\Sigma\mid W)}{\sup_{\alpha^{\prime}\in A,\,\beta^{\prime}\in\mathbb{R},\,\Sigma^{\prime}\in\mathbb{S}^{2}_{+}}f_{\mathrm{S}}(\alpha^{\prime},\beta^{\prime},\Sigma^{\prime}\mid W)}.$ (6)
To solve the optimisation problem in (6), we marginalise out $\Sigma$ first to simplify the problem. To marginalise out $\Sigma$, we plug in the maximum-likelihood estimator of the reduced-form covariance $\Psi$ given by
@@
$R(\beta)\hat{\Sigma}(\beta)R(\beta)^{\intercal}=\hat{\Psi}$; (7)
indeed, for any $\beta$, we can ensure that (7) holds by taking
@@
where the “*” is expressing symmetry. This ensures that whatever the considered value of $\beta$, we can still achieve the global maximum in $\Psi$. Fixing $\Sigma=\hat{\Sigma}(\beta)$, we obtain from equation (5)
@@
The aim is to maximise $f_{\mathrm{S}}(\alpha,\beta\mid W)$, or equivalently $\log f_{\mathrm{S}}(\alpha,\beta\mid W)$, with respect to $\alpha$ given the constraint $\alpha\in A$. For a fixed $\beta$, it is sufficient to minimise $(\alpha-t(\beta))^{\intercal}(Z^{\intercal}Z)(\alpha-t(\beta))$ under the considered constraint $\alpha\in A$. Clearly, $t(\beta)$ is the minimiser whenever $t(\beta)\in A$. Otherwise, the minimiser is the point $\alpha^{*}\in A$ that is closest to $t(\beta)$ in the distance implied by $Z^{\intercal}Z$, i.e. the projection of $t(\beta)$ onto $A$ with respect to the metric induced by $Z^{\intercal}Z$. Equivalently, the solution is the point that minimises the Mahalanobis distance from a distribution with mean vector $t$ and covariance matrix $\hat{\sigma}_{11}(\beta)(Z^{\intercal}Z)^{-1}$.
When $A$ is a rectangle, i.e. of the form $A=[a_{1},b_{1}]\times\ldots\times[a_{p},b_{p}]$, this is a standard quadratic programming problem, which we solve using the JuMP.jl package (lubin_jump_2023). If $Z^{\intercal}Z$ is diagonal, this simplifies to clipping component-wise.
Alternatively, we can also bound a norm of $\alpha$, that is specify the constraint set as $A_{\tau}=\left\{\alpha:\lVert\alpha\rVert\leq\tau\right\}$, where the threshold $\tau$ is the maximum invalidity budget across all instruments (similar to penn_budgetiv_2025). This may be easier to specify in some settings. For the $\ell_{2}$ norm, this turns the optimisation problem into a Ridge-type regression problem, and we have that the optimal $\alpha$ is
@@
where $\lambda$ is the Lagrange multiplier corresponding to a specific threshold $\tau$. In practice, one has to find $\lambda$ such that $\lVert\alpha\rVert_{2}^{2}=\tau$. For the $\ell_{1}$ norm, the optimisation problem could be solved by a LARS-type (efron_least_2004) algorithm, but we leave details for future work.
@@ A.3 Proof of Proposition 2
Let $\alpha_{0}\in\mathbb{R}^{p}$ denote the true data-generating value of $\alpha$. Then, the validified posterior possibility as a functional of $W$, $\pi_{W}(\beta\mid\{\alpha_{0}\})$, is a probability integral transform, and its distribution is stochastically greater or equal to a Uniform distribution on $[0,1]$ (see for example casella_statistical_2002, Section 2.1). Thus, for any $\delta\in[0,1]$, we have $P_{\beta}\left(\pi_{W}(\beta\mid\{\alpha_{0}\})\leq\delta\right)\leq\delta$. For any violation set that contains the true value, $A\supseteq\{\alpha_{0}\}$, the validified posterior becomes no more informative, i.e., $\pi_{w}(\beta\mid A)\geq\pi_{w}(\beta\mid\{\alpha_{0}\})$, such that
@@
Appendix B Additional details on the experiments
Here, we provide additional details on the implementation of the competing methods in our simulation experiments. For the plausible GMM (PGMM) estimator, we put a baseline Gaussian prior with identity prior covariance on the moment restriction. This puts all values of $\alpha$ considered (transformed to the moment restriction) well within the centre of that distribution. To implement BudgetIV, we use the code available at https://github.com/jpenn2023/budgetIVr. Their method has two hyperparameters $b$ and $\tau$, which jointly specify how many components of $\alpha$ can exceed the threshold $\tau$. We set $b=1$ and specify the budget $\tau$ analogously to the thresholds we use for our possibilistic approach. This means that at least one instrument has to be valid in the sense that $\lvert\alpha_{i}\rvert\leq\tau$. In the multiple instrument case, we would rather set $b=p=5$, but this never returns a plausible set. This explains why BudgetIV is very conservative in the multiple instrument case. The CIIV results are based on the implementation available at https://github.com/xlbristol/CIIV with default settings. The gIVBMA method is implemented by https://github.com/gregorsteiner/gIVBMA.jl and we choose the hyper-$g/n$ prior specification.

### v2

@@ 5 Conclusion
In this paper, we propose a method for instrumental variable regression based on possibility theory. This allows for valid inference under some user-specified violations of the instrument exogeneity assumption. The proposed method also provides a very natural way to perform sensitivity analyses. The resulting inference is based on first principles and directly reflects the epistemic uncertainty about the parameters. Our approach performs well in simulation experiments and delivers credible results in real-data applications.
@@ A.1 The structural posterior possibility function
$\begin{aligned} \log f_{\mathrm{S}}(\alpha,\beta,\Sigma\mid W)&=-\frac{n}{2}\log\lvert R(\beta)\Sigma R(\beta)^{\intercal}\rvert\\ &-\frac{1}{2}\text{tr}\left((R(\beta)\Sigma R(\beta)^{\intercal})^{-1}W^{\intercal}M_{Z}W\right)\\ &-\frac{1}{2}(\alpha-t(\beta))^{\intercal}\frac{Z^{\intercal}Z}{\sigma_{11}}(\alpha-t(\beta)).\end{aligned}$ (A.1)
@@ A.2 Proof of Proposition 1 and computational considerations
$f(\beta\mid\alpha\in A,W)=\sup_{\alpha\in A,\,\Sigma\in\mathbb{S}^{2}_{+}}\frac{f_{\mathrm{S}}(\alpha,\beta,\Sigma\mid W)}{\sup_{\alpha^{\prime}\in A,\,\beta^{\prime}\in\mathbb{R},\,\Sigma^{\prime}\in\mathbb{S}^{2}_{+}}f_{\mathrm{S}}(\alpha^{\prime},\beta^{\prime},\Sigma^{\prime}\mid W)}.$ (A.2)
To solve the optimisation problem in (A.2), we marginalise out $\Sigma$ first to simplify the problem. To marginalise out $\Sigma$, we plug in the maximum-likelihood estimator of the reduced-form covariance $\Psi$ given by
@@
$R(\beta)\hat{\Sigma}(\beta)R(\beta)^{\intercal}=\hat{\Psi}$; (A.3)
indeed, for any $\beta$, we can ensure that (A.3) holds by taking
@@
where the “*” is expressing symmetry. This ensures that whatever the considered value of $\beta$, we can still achieve the global maximum in $\Psi$. Fixing $\Sigma=\hat{\Sigma}(\beta)$, we obtain from equation (A.1)
@@
The aim is to maximise $f_{\mathrm{S}}(\alpha,\beta\mid W)$, or equivalently $\log f_{\mathrm{S}}(\alpha,\beta\mid W)$, with respect to $\alpha$ given the constraint $\alpha\in A$. For a fixed $\beta$, it is sufficient to minimise $(\alpha-t(\beta))^{\intercal}(Z^{\intercal}Z)(\alpha-t(\beta))$ under the considered constraint $\alpha\in A$. Clearly, $t(\beta)$ is the minimiser whenever $t(\beta)\in A$. Otherwise, the minimiser is the point $\alpha^{*}\in A$ that is closest to $t(\beta)$ in the distance implied by $Z^{\intercal}Z$, i.e. the projection of $t(\beta)$ onto $A$ with respect to the metric induced by $Z^{\intercal}Z$. Equivalently, the solution is the point that minimises the Mahalanobis distance from a distribution with mean vector $t$ and covariance matrix $\hat{\sigma}_{11}(\beta)(Z^{\intercal}Z)^{-1}$. Figure A.1 illustrates this in two dimensions.
[…]
Figure A.1: A geometric illustration of our method for $p=2$: The causal effect $\beta$ is partially identified where the affine subspace $t(\beta)=\hat{\gamma}_{1}-\beta\hat{\gamma}_{2}$ intersects the tolerated region AA. At these values of $\beta$, the corresponding $\alpha$ is precisely $t(\beta)$, and therefore, the conditional possibility is $1$. For all other values of $\beta$, the optimal $\alpha$ is the projection onto AA w.r.t. the metric induced by $Z^{\intercal}Z$.
When $A$ is a rectangle, i.e. of the form $A=[a_{1},b_{1}]\times\ldots\times[a_{p},b_{p}]$, this is a standard quadratic programming problem, which we solve using the JuMP.jl package (Lubin et al.,, 2023). If $Z^{\intercal}Z$ is diagonal, this simplifies to clipping component-wise.
Alternatively, we can also bound a norm of $\alpha$, that is specify the constraint set as $A_{\tau}=\left\{\alpha:\lVert\alpha\rVert\leq\tau\right\}$, where the threshold $\tau$ is the maximum invalidity budget across all instruments (similar to Penn et al.,, 2025). This may be easier to specify in some settings. For the $\ell_{2}$ norm, this turns the optimisation problem into a Ridge-type regression problem, and we have that the optimal $\alpha$ is
@@
where $\lambda$ is the Lagrange multiplier corresponding to a specific threshold $\tau$. In practice, one has to find $\lambda$ such that $\lVert\alpha\rVert_{2}^{2}=\tau$. For the $\ell_{1}$ norm, the optimisation problem could be solved by a LARS-type (Efron et al.,, 2004) algorithm, but we leave details for future work.
@@ A.3 Proof of Proposition 2
Let $\alpha_{0}\in\mathbb{R}^{p}$ denote the true data-generating value of $\alpha$. Then, the validified posterior possibility as a functional of $W$, $\pi_{W}(\beta\mid\{\alpha_{0}\})$, is a probability integral transform, and its distribution is stochastically greater or equal to a Uniform distribution on $[0,1]$ (see for example Casella and Berger,, 2002, Section 2.1). Thus, for any $\delta\in[0,1]$, we have $P_{\beta}\left(\pi_{W}(\beta\mid\{\alpha_{0}\})\leq\delta\right)\leq\delta$. For any violation set that contains the true value, $A\supseteq\{\alpha_{0}\}$, the validified posterior becomes no more informative, i.e., $\pi_{w}(\beta\mid A)\geq\pi_{w}(\beta\mid\{\alpha_{0}\})$, such that
@@
Appendix B Additional details on the simulation experiments
Here, we provide additional details on the implementation of the competing methods in our simulation experiments. For the plausible GMM (PGMM) estimator, we put a baseline Gaussian prior with identity prior covariance on the moment restriction. This puts all values of $\alpha$ considered (transformed to the moment restriction) well within the centre of that distribution. To implement BudgetIV, we use the code available at https://github.com/jpenn2023/budgetIVr. Their method has two hyperparameters, $b$ and $\tau$, which specify that at least $b$ components of $\alpha$ cannot exceed the threshold $\tau$. We set $b=1$ and select the budget $\tau$ analogously to the thresholds we use for our possibilistic approach. This means that at least one instrument has to correspond to $\lvert\alpha_{i}\rvert\leq\tau$. In the case of multiple instruments, setting $b=p=5$ would provide a fairer comparison with our approach, since then all instruments are restricted to be potentially invalid only up to degree $\tau$. However, this configuration rarely yields a feasible set for the treatment effect, i.e., one where the affine subspace intersects the partial identification region, in which case BudgetIV returns no set. In contrast, our approach still produces a posterior possibility function based on projection, where the mode corresponds to the value of $\beta$ that implies an $\alpha$ closest to the violation set $A$. For this reason, we maintain $b=1$ for BudgetIV, which then always returns a plausible set. Unsurprisingly, this additional flexibility renders BudgetIV very conservative in the multiple instrument case. The CIIV results are based on the implementation available at https://github.com/xlbristol/CIIV with default settings. The gIVBMA method is implemented by https://github.com/gregorsteiner/gIVBMA.jl and we choose the hyper-$g/n$ prior specification.
Appendix C Additional details on the real-data examples
In the context of the growth example, Table C.1 presents $95\%$ uncertainty intervals for our approach along with those obtained by standard TSLS or different variants of PGMM. Table C.2 displays upper and lower probabilities for the hypothesis $\beta>0$. These are discussed in more detail in the main text.
For the returns to schooling example, Table C.3 presents upper and lower probabilites for the hypothesis of positive returns to schooling.
Table C.1: The effect of institutions on economic growth: $95\%$ uncertainty intervals for possibilistic IV, TSLS, and PGMM (taken from Chernozhukov et al.,, 2025). The intervals for our approach are based on the $\chi^{2}$ approximation. PGMM-u, PGMM-g and PGMM(d)-g refer to PGMM with, respectively, uniform prior, baseline Gaussian prior and diffuse Gaussian prior.
Method $95\%$ Interval
Possibilistic IV ($A=\{0\}$) $[0.69,1.62]$
Possibilistic IV (A=[0.1,0.1]A=[-0.1,0.1]) $[0.53,1.96]$
TSLS $[0.56,1.38]$
PGMM-u $[0.58,3.65]$
PGMM-g $[0.49,3.79]$
PGMM(d)-g $[0.22,3.81]$
Table C.2: The effect of institutions on economic growth: Lower and upper probabilities for the hypothesis $\beta>0$ under the constraint $\alpha\in A$ (based on the MC approximation with $M=10,000$).
A Lower Upper
$\{0\}$ 1 1
[0.1,0.1][-0.1,0.1] 1 1
[0.2,0.2][-0.2,0.2] 1 1
[0.3,0.3][-0.3,0.3] 1 1
[0.4,0.4][-0.4,0.4] 0.927 1
[0.5,0.5][-0.5,0.5] 0.603 1
Table C.3: The returns to schooling: Lower and upper probabilities for the hypothesis $\beta>0$ under the constraint $\alpha\in A$ (based on the MC approximation with $M=5,000$).
A Lower Upper
$\{0\}$ 1 1
$[0,0.01]$ 1 1
$[0,0.02]$ 0.887 1
$[0,0.03]$ 0.748 1
$[0,0.04]$ 0.517 1